i need to find the diaganol of my computer monitor using a^2+b^2=c^2
L=13"
W=8"
D=????

Found 2 solutions by oscargut, checkley71:
Answer by oscargut(2103)

D^2=13^2+8^2=169+64=233 then D=sqrt(233)"

Answer by checkley71(8403)

13^2+8^2=c^2
169+64=c^2
233=c^2
c=sqrt233
c=15.26"